# Reassessing constant-roll Warm Inflation

### v1

@@ I Introduction
On the one hand, cosmic inflation Kazanas:1980tx ; Guth:1980zm ; Sato:1981ds ; Sato:1980yn ; Linde:1981mu ; Albrecht:1982wi solves the fine-tuning problems of the Big Bang cosmology, i.e. the horizon problem, the flatness problem and the monopole problem. On the other hand, it helps generate inhomogeneties in an otherwise homogeneous and isotropic Friedmann-Lemaitre-Robertson-Walker (FLRW) background, which eventually give rise to the large scale structures we see around us today as well as the fluctuations in the temperature of the Cosmic Microwave Background Radiation (CMBR) which have been measured to a very good accuracy in recent times. The dynamics of the inflationary phase is often portrayed as the dynamics of a single scalar field which is slowly rolling down its approximately flat potential. The slow-roll dynamics of the inflaton field is often invoked in any standard inflationary dynamics as it helps gracefully exit the inflationary phase as well as generate nearly scale-invariant spectra of primordial perturbations (both scalar and tensor), and thus makes definite predictions which can be tested through the observations of the CMBR and the large scale structures Riotto:2002yw ; Baumann:2009ds ; Mishra:2024axb . This simple vanilla model of single field slow-roll inflation is in good agreement with the current data Planck:2018jri . Significant departure from the slow-roll dynamics, as it happens during an ultra-slow-roll phase Dimopoulos:2017ged for example, results in departure from the nearly scale-invariant spectra that slow-roll generates. It has been shown that a transient phase of Ultra-slow-roll yields large primordial perturbations which eventually collapse to form Primordial Black Holes Motohashi:2017kbs , a coveted candidate for Dark Matter Carr:2016drx .
@@
where the constant-roll parameter $\beta$ is a constant and can take both positive and negative values. Note that the constant-roll parameter $n$ in Martin:2012pe is equal to $-3\beta$ in our case. The above equation can be rewritten as $d\ln\dot{\phi}/dN=-3\beta$, where $N$ is the number of $e-$folds. This clearly shows that $\dot{\phi}$ changes with a constant rate ($-3\beta$) with $e-$folds, and hence the name. If $|\beta|\ll 1$, then one can ignore $\ddot{\phi}$ with respect to the $3H\dot{\phi}$ term, and in such cases the dynamics boils down to the familiar slow-roll dynamics. On the other hand, $\beta=1$ corresponds to the ultra-slow-roll dynamics for a phase when $\partial V/\partial\phi=0$, where $V$ represents the inflaton potential Dimopoulos:2017ged . One can get an estimate of the allowed values of $\beta$ in two different ways. First of all, it has been shown in Morse:2018kda ; Lin:2019fcz that only small-$\beta$ constant-roll models ($\beta<0.5$) can be universal attractors, whereas large-$\beta$ constant-roll ($\beta>0.5$) should be treated as an initial transient phase followed by an evolution dominated by slow-roll dynamics. The constant-roll parameter $\bar{\eta}$ in Morse:2018kda ; Lin:2019fcz is equal to $3\beta$ in our case. Secondly, the perturbations generated during a constant-roll inflationary phase (not the initial transient phase followed by slow-roll dynamics) are compatible with current data only when $\beta$ is negative and of the order $10^{-3}$ Motohashi:2017aob . Note that the constant roll parameter $\beta$ in Motohashi:2017aob is equal to $-3\beta$ in our case. However, it is shown in Gao:2019sbz that $\beta\sim 1$ is also allowed observationally, but such constant-roll inflation models fails to gracefully exit an inflationary phase. Note that the constant-roll parameter $\eta_{H}$ in Gao:2019sbz is equal to $3\beta$ in our case. This indicates that despite imposing the constant-roll condition, given in Eq. (1), on a single-field inflationary dynamics, to be an attractor solution as well as to remain viable by the data (in addition, to be able to gracefully exit an inflationary phase), the dynamics should not differ much from a slow-roll evolution.
@@
In this article we will reassess the constant-roll dynamics in a variant inflationary scenario, known as Warm Inflation Berera:1995ie . Unlike in the standard single-field inflationary scenario described above, where the couplings of the inflaton field is generally ignored during an inflationary phase, in Warm Inflation (WI) these couplings play a significant role in dissipating inflaton’s energy to maintain a constant radiation bath throughout inflation (for more recent reviews on WI, see Kamali:2023lzq ; Berera:2023liv ). In doing so, WI can gracefully exit to a radiation dominated era, whereas the more standard inflationary scenario, which we will refer to as Cold Inflation (CI) henceforth, requires a subsequent reheating phase, dynamics of which is still largely unknown. Also, the requirement of reheating post inflation in CI scenarios restricts the classes of inflationary potentials to the ones with a minimum. Otherwise, one requires other mechanism to reheat the universe, such as gravitational reheating Ford:1986sy ; Chun:2009yu for example, apart from standard reheating dynamics via the inflaton’s oscillation at the minima of the potential Mishra:2024axb . As WI doesn’t require a subsequent reheating phase, such scenarios can accommodate larger classes of inflaton potentials. For example, generalised runaway potentials, which do not have minima, have been explored in Das:2020xmh and was shown that the WI dynamics smoothly transits to a radiation dominated era. However, graceful exit in WI is a much more complex process than in CI as has been pointed out in Das:2020lut . Some of the attractive features of WI include (a) generation of smaller tensor-to-scalar ratio which helps accommodate potentials, like quartic and quadratic, which are otherwise ruled out in CI for generating way too much tensor perturbations compared to scalar perturbations Bartrum:2013fia , (b) generation of Primordial Black Holes without departing from the slow-roll dynamics Arya:2019wck ; Bastero-Gil:2021fac ; Correa:2022ngq ; Arya:2023pod , (c) accommodating the de Sitter Swampland Conjecture Ooguri:2018wrx ; Garg:2018reu naturally Das:2018hqy ; Motaharfar:2018zyb ; Das:2018rpg ; Das:2019hto ; Das:2019acf which otherwise disfavours the standard CI dynamics Kinney:2018nny , to name of few.
@@ IV.4 The case with $\Upsilon$ as a function of $\phi$
where the power of $T$ and $\phi$, i.e. $p$ and $c$ respectively, can take integer values (positive or negative), $C_{\Upsilon}$ is a dimensionless quantity which depends on the microscopic details of the WI model under consideration and $M$ is some appropriate mass scale so that the dimensionality of $\Upsilon$ is kept at $[\Upsilon]=[{\rm mass}]$.
@@ Acknowledgements.
The work of S.D. is supported by the Start-up Research Grant (SRG) awarded by Science and Engineering Research Board (SERB) [Grant No: SRG/2023/000101/PMS]. S.D. is also thankful to Axis Bank and acknowledges the financial support obtained from them which partially supports this research. S.D. thanks Rudnei Ramos for many useful discussions on Warm Inflation on various occasions.

### v2

@@ I Introduction
On the one hand, cosmic inflation Kazanas:1980tx ; Guth:1980zm ; Sato:1981ds ; Sato:1980yn ; Linde:1981mu ; Albrecht:1982wi solves the fine-tuning problems of the Big Bang cosmology, i.e. the horizon problem, the flatness problem and the monopole problem. On the other hand, it helps generate inhomogeneties in an otherwise homogeneous and isotropic Friedmann-Lemaitre-Robertson-Walker (FLRW) background, which eventually give rise to the large scale structures we see around us today as well as the fluctuations in the temperature of the Cosmic Microwave Background Radiation (CMBR) which have been measured to a very good accuracy in recent times. The dynamics of the inflationary phase is often portrayed as the dynamics of a single scalar field which is slowly rolling down its approximately flat potential. The slow-roll dynamics of the inflaton field is often invoked in any standard inflationary dynamics as it helps gracefully exit the inflationary phase as well as generate nearly scale-invariant spectra of primordial perturbations (both scalar and tensor), and thus makes definite predictions which can be tested through the observations of the CMBR and the large scale structures Riotto:2002yw ; Baumann:2009ds ; Mishra:2024axb . This simple vanilla model of single field slow-roll inflation is in good agreement with the current data Planck:2018jri . Significant departure from the slow-roll dynamics, as it happens during an ultra-slow-roll phase Kinney:2005vj ; Dimopoulos:2017ged for example, results in departure from the nearly scale-invariant spectra that slow-roll generates. It has been shown that a transient phase of Ultra-slow-roll yields large primordial perturbations which eventually collapse to form Primordial Black Holes Motohashi:2017kbs , a coveted candidate for Dark Matter Carr:2016drx .
@@
where the constant-roll parameter $\beta$ is a constant and can take both positive and negative values. Note that the constant-roll parameter $n$ in Martin:2012pe is equal to $-3\beta$ in our case. The above equation can be rewritten as $d\ln\dot{\phi}/dN=-3\beta$, where $N$ is the number of $e-$folds. This clearly shows that $\dot{\phi}$ changes with a constant rate ($-3\beta$) with $e-$folds, and hence the name. If $|\beta|\ll 1$, then one can ignore $\ddot{\phi}$ with respect to the $3H\dot{\phi}$ term, and in such cases the dynamics boils down to the familiar slow-roll dynamics. On the other hand, $\beta=1$ corresponds to the ultra-slow-roll dynamics for a phase when $\partial V/\partial\phi=0$, where $V$ represents the inflaton potential Kinney:2005vj ; Dimopoulos:2017ged . One can get an estimate of the allowed values of $\beta$ in two different ways. First of all, it has been shown in Morse:2018kda ; Lin:2019fcz that only small-$\beta$ constant-roll models ($\beta<0.5$) can be universal attractors, whereas large-$\beta$ constant-roll ($\beta>0.5$) should be treated as an initial transient phase followed by an evolution dominated by slow-roll dynamics. The constant-roll parameter $\bar{\eta}$ in Morse:2018kda ; Lin:2019fcz is equal to $3\beta$ in our case. Secondly, the perturbations generated during a constant-roll inflationary phase (not the initial transient phase followed by slow-roll dynamics) are compatible with current data only when $\beta$ is negative and of the order $10^{-3}$ Motohashi:2017aob . Note that the constant roll parameter $\beta$ in Motohashi:2017aob is equal to $-3\beta$ in our case. However, it is shown in Gao:2019sbz that $\beta\sim 1$ is also allowed observationally, but such constant-roll inflation models fails to gracefully exit an inflationary phase. Note that the constant-roll parameter $\eta_{H}$ in Gao:2019sbz is equal to $3\beta$ in our case. This indicates that despite imposing the constant-roll condition, given in Eq. (1), on a single-field inflationary dynamics, to be an attractor solution as well as to remain viable by the data (in addition, to be able to gracefully exit an inflationary phase), the dynamics should not differ much from a slow-roll evolution.
@@
In this article we will reassess the constant-roll dynamics in a variant inflationary scenario, known as Warm Inflation Berera:1995ie . Unlike in the standard single-field inflationary scenario described above, where the couplings of the inflaton field is generally ignored during an inflationary phase, in Warm Inflation (WI) these couplings play a significant role in dissipating inflaton’s energy to maintain a constant radiation bath throughout inflation (for more recent reviews on WI, see Kamali:2023lzq ; Berera:2023liv ). In doing so, WI can gracefully exit to a radiation dominated era, whereas the more standard inflationary scenario, which we will refer to as Cold Inflation (CI) henceforth, requires a subsequent reheating phase, dynamics of which is still largely unknown. Also, the requirement of reheating post inflation in CI scenarios restricts the classes of inflationary potentials to the ones with a minimum. Otherwise, one requires other mechanism to reheat the universe, such as gravitational reheating Ford:1986sy ; Chun:2009yu for example, apart from standard reheating dynamics via the inflaton’s oscillation at the minima of the potential Mishra:2024axb . As WI doesn’t require a subsequent reheating phase, such scenarios can accommodate larger classes of inflaton potentials. For example, generalised runaway potentials, which do not have minima, have been explored in Das:2020xmh and was shown that the WI dynamics smoothly transits to a radiation dominated era. However, graceful exit in WI is a much more complex process than in CI as has been pointed out in Das:2020lut . Some of the attractive features of WI include (a) generation of smaller tensor-to-scalar ratio which helps accommodate potentials, like quartic and quadratic, which are otherwise ruled out in CI for generating way too much tensor perturbations compared to scalar perturbations Bartrum:2013fia ,  [^1] Recently two numerical codes have been developed to compute perturbations in WI Montefalcone:2023pvh ; Ballesteros:2023dno . (b) generation of Primordial Black Holes without departing from the slow-roll dynamics Arya:2019wck ; Bastero-Gil:2021fac ; Correa:2022ngq ; Arya:2023pod , (c) accommodating the de Sitter Swampland Conjecture Ooguri:2018wrx ; Garg:2018reu naturally Das:2018hqy ; Motaharfar:2018zyb ; Das:2018rpg ; Das:2019hto ; Das:2019acf which otherwise disfavours the standard CI dynamics Kinney:2018nny , to name of few.
@@ II A fresh look at constant-roll Cold Inflation: Graceful exit condition
Note that as this equation is derived from Hubble slow-roll parameter which depends only on the background evolution (and not on the form of the inflaton potential), this equation is valid for slow-roll or beyond slow-roll CI or WI.
@@ IV.4 The case with $\Upsilon$ as a function of $\phi$
where the power of $T$ and $\phi$, i.e. $p$ and $c$ respectively, can take integer values (positive or negative), $C_{\Upsilon}$ is a dimensionless quantity which depends on the microscopic details of the WI model under consideration and $M$ is some appropriate mass scale so that the dimensionality of $\Upsilon$ is kept at $[\Upsilon]=[{\rm mass}]$. The stability of WI models restricts the dependance of $T$ of $\Upsilon$ to $|p|\leq 4$ Moss:2008yb .
@@ Acknowledgements.
The work of S.D. is supported by the Start-up Research Grant (SRG) awarded by Science and Engineering Research Board (SERB) [Grant No: SRG/2023/000101/PMS], Department of Science and Technology (DST), Government of India. S.D. is also thankful to Axis Bank and acknowledges the financial support obtained from them which partially supports this research. S.D. thanks Rudnei Ramos for many useful discussions on Warm Inflation on various occasions.